# Split Gibbs Discrete Diffusion Posterior Sampling

### v2

@@ Abstract
We study the problem of posterior sampling in discrete-state spaces using discrete diffusion models. While posterior sampling methods for continuous diffusion models have achieved remarkable progress, analogous methods for discrete diffusion models remain challenging. In this work, we introduce a principled plug-and-play discrete diffusion posterior sampling algorithm based on split Gibbs sampling, which we call SGDD. Our algorithm enables reward-guided generation and solving inverse problems in discrete-state spaces. We demonstrate the convergence of SGDD to the target posterior distribution and verify this through controlled experiments on synthetic benchmarks. Our method enjoys state-of-the-art posterior sampling performance on a range of benchmarks for discrete data, including DNA sequence design, discrete image inverse problems, and music infilling, achieving more than $30\%$ improved performance compared to existing baselines. Our code is available at SGDD.
@@ 1 Introduction
We validate SGDD on diverse inverse problems and reward-guided generation problems that involve discrete data. We demonstrate that SGDD converges to the true posterior distribution on synthetic data, that SGDD solves discrete inverse problems accurately and efficiently on discretized image data and monophonic music data, and that SGDD excels at reward-guided sampling on guided DNA generation. Our method demonstrates strong sampling quality in all experiments. For instance, we achieve a $42\%$ higher median activity in enhancer DNA design, an $8.36\ \mathrm{dB}$ improvement of PSNR values in solving the XOR inverse problem on MNIST, and over 2×2$\times$2 × smaller Hellinger distance in music infilling, compared to existing methods.
@@ 2.2 Diffusion Posterior Sampling Methods
Existing diffusion posterior sampling methods leverage (continuous) diffusion models as plug-and-play priors and modify Eq. 1 to generate samples from the posterior distribution $p({\mathbf{x}}|{\mathbf{y}})\ \propto\ p({\mathbf{x}})p({\mathbf{y}}|{\mathbf{x}})$. These methods can be broadly classified into four categories [9, 53].
Guidance-based methods [7, 36, 35, 24, 46, 52] perform posterior sampling by adding an additional likelihood score term $\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{y}}|{\mathbf{x}}_{t})$ to the score function in Eq. 1, and modifies the reverse SDE as: $\mathrm{d}{\mathbf{x}}_{t}=-2\dot{\sigma}_{t}\sigma_{t}(\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{x}}_{t};\sigma_{t})+\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{y}}|{\mathbf{x}}_{t}))\mathrm{d}t+\sqrt{2\dot{\sigma}_{t}\sigma_{t}}\mathrm{d}{\mathbf{w}}_{t}.$ The intractable likelihood score term is approximated with various assumptions. However, these methods cannot be directly applied to discrete posterior sampling problems, where $\log p({\mathbf{y}}|{\mathbf{x}})$ is defined only on a finite support and thus does not have such gradient information.
Sequential Monte Carlo (SMC) methods sample batches of particles iteratively from a series of distributions, which converge to the posterior distribution in the limit. This class of methods [5, 12, 26, 47] exploits the sequential diffusion sampling process to sample from the posterior distribution. Similar ideas have been applied in Li et al. [27] for reward-guided generation using discrete diffusion models, which rely on approximating a soft value function at each iteration.
@@
Variable splitting methods [48, 51, 54, 34, 49] decompose posterior sampling into two simpler alternating steps. A representative of this category is the split Gibbs sampler (see Section 3), where the variable ${\mathbf{x}}_{t}$ is split into two random variables that both eventually converge to the posterior distribution. These methods are primarily designed for continuous diffusion models with Gaussian kernels, leveraging Langevin dynamics or optimization in Euclidean space as subroutines. Murata et al. [31] extends this category to VQ-diffusion models in the discrete latent space using the Gumbel-softmax reparameterization trick. However, this approach is limited to discrete latent spaces with an underlying continuous embedding due to its reliance on softmax dequantization. In this work, we generalize the split Gibbs sampler to posterior sampling in categorical data by further exploiting the unique properties of discrete diffusion models.
@@ 3.1 Split Gibbs for plug-and-play posterior sampling
where $D({\mathbf{x}},{\mathbf{z}};\eta)$ measures the distance between ${\mathbf{x}}$ and ${\mathbf{z}}$, and $\eta>0$ is a parameter that controls the strength of regularization. While prior works [8, 48, 49] consider $D({\mathbf{x}},{\mathbf{z}};\eta)=\frac{\|{\mathbf{x}}-{\mathbf{z}}\|^{2}_{2}}{2\eta^{2}}$, the potential function can be generalized to any continuous function $D$, such that $D({\mathbf{x}},{\mathbf{z}};\eta)\to\infty$ as $\eta\to 0$ for any ${\mathbf{x}}\neq{\mathbf{z}}$. As shown in Section A.3, this ensures that both marginal distributions,
$\pi^{X}({\mathbf{x}};\eta):=\int\pi({\mathbf{x}},{\mathbf{z}};\eta)\mathrm{d}{\mathbf{z}},\ \ \text{and}\ \ \pi^{Z}({\mathbf{z}};\eta):=\int\pi({\mathbf{x}};{\mathbf{z}};\eta)\mathrm{d}{\mathbf{x}}$
converges to the posterior distribution $p({\mathbf{x}}|{\mathbf{y}})$. The decoupling of the prior and likelihood in Eq. 7 enables split Gibbs sampling, which alternates between two steps:
@@ 2. Prior sampling step:
Prior work has studied split Gibbs posterior sampling for continuous diffusion models [8, 48, 49]. These methods specify the regularization potential as $D({\mathbf{x}},{\mathbf{z}};\eta)=\frac{\|{\mathbf{x}}-{\mathbf{z}}\|_{2}^{2}}{2\eta^{2}}$, which transforms Eq. 9 into a Gaussian denoising problem solvable by a diffusion model. A key challenge, which we tackle, is developing a performant approach for the discrete diffusion setting that is easy to implement, and enjoys rigorous guarantees (i.e., sampling from the correct posterior distribution).
@@ 3.2 Prior Step with Discrete Diffusion Models
Suppose $p({\mathbf{x}})$ is a discrete-state distribution over $\mathcal{X}^{D}$ modeled by a diffusion process. We consider a discrete diffusion model with uniform transition kernel ${\bm{Q}}_{t}^{\mathsf{fw}}=\frac{1}{N}\mathbf{1}\mathbf{1}^{T}-{\bm{I}}$. To connect split Gibbs sampling to discrete diffusion models, we specify the potential function $D({\mathbf{x}};{\mathbf{z}};\eta)$ as:
@@
where $d({\mathbf{x}},{\mathbf{z}})$ denotes the Hamming distance between ${\mathbf{x}}$ and ${\mathbf{z}}$. When $\eta\to 0^{+}$, the regularization potential $D({\mathbf{x}},{\mathbf{z}};\eta)$ goes to infinity unless $d({\mathbf{x}},{\mathbf{z}})=0$, ensuring the convergence of marginal distributions to $p({\mathbf{x}}|{\mathbf{y}})$. Given Eq. 10, the prior step from Eq. 9 can be written as
@@
where $\beta_{t}=\frac{N-1}{N}(1-e^{-\sigma_{t}})$. Thus, sampling from Eq. 11 is equivalent to unconditional generation from $p({\mathbf{x}}_{0}|{\mathbf{x}}_{t})$ when $\tilde{\beta}=\beta_{t}$, i.e., $\eta=\sigma_{t}$. Therefore, we can solve the prior sampling problem by simulating a partial discrete diffusion sampler starting from $\sigma_{t}=\eta$ and ${\mathbf{x}}_{t}={\mathbf{z}}^{(k)}$.
@@ 3.3 Likelihood Sampling Step
With the potential function $D({\mathbf{x}},{\mathbf{z}};\eta)$ specified in Section 3.2, at iteration $k$, the likelihood sampling step from Eq. 8 can be written as:
@@
Since the unnormalized probability density function of $\pi({\mathbf{x}}={\mathbf{x}}^{(k)},{\mathbf{z}};\eta)$ is accessible, we can efficiently sample from Eq. 13 using Markov Chain Monte Carlo methods.
@@
     Likelihood step following Eq. 13: ${\mathbf{z}}^{(k)}\sim\pi({\mathbf{x}}={\mathbf{x}}^{(k)},{\mathbf{z}};\eta_{k}).$
     Prior step following Eq. 11: ${\mathbf{x}}^{(k+1)}\sim\pi({\mathbf{x}},{\mathbf{z}}={\mathbf{z}}^{(k)};\eta_{k})$
@@ 3.4 Overall Algorithm
We now summarize the complete SGDD algorithm. Like in Wu et al. [48], our algorithm alternates between likelihood steps and prior steps while employing an annealing schedule $\{\eta_{k}\}$, which starts at a large $\eta_{0}$ and gradually decays to $\eta_{K}\to 0$. This annealing scheme accelerates the mixing time of the Markov chain, which we further provide some theoretical intuition in Section B.1; and ensures the convergence of $\pi^{X}({\mathbf{x}};\eta)$ and $\pi^{Z}({\mathbf{z}};\eta)$ to $p({\mathbf{x}}|{\mathbf{y}})$ as $\eta\to 0$. We present the complete pseudocode of our method in Algorithm 1 and further explain the choice of the noise schedule in Section B.1.
@@ 3.5 Convergence of Split Gibbs Diffusion Posterior Sampling
$\frac{1}{K}\sum_{k=0}^{K-1}\underbrace{\frac{1}{t^{*}}\int_{k(t^{*}+1)+1}^{(k+1)(t^{*}+1)}\mathsf{FI}_{{\bm{Q}}_{t}}(\mu_{t}\|\pi_{t})}_{\textit{average Fisher divergence in the k-th prior step}}\leq\underbrace{\frac{2\mathsf{KL}(\mu_{0}\|\pi_{0})}{Kt^{*}}}_{\textit{convergence by}\ O(1/K)}+\underbrace{\frac{4M\epsilon}{c}}_{\textit{score error}}+\underbrace{\frac{2MLt^{*}}{cH}}_{\textit{discretization error}}.$ (14)
where $\|\frac{s_{\theta}(\cdot,t)-s(\cdot,t)}{s(\cdot,t)}\|_{\infty}\leq\epsilon<1$, and $L,M,c$ are positive constants defined in Appendix A.
We include the proof in Appendix A. Theorem 1 states that the average Fisher divergence of the non-stationary process to the stationary process in all prior steps converges at the rate of $O(1/K)$, up to a constant error term. This result extends our theoretical understanding of diffusion posterior sampling by generalizing the analysis [41, 48] on SDEs to general Markov processes using the free-energy-rate-functional-relative-Fisher (FIR) inequality [16]. Moreover, compared to existing analyses for continuous diffusion models, our analysis accounts not only for the imperfect score function but also for the discretization error in solving CTMCs as well.
Unlike SMC-based methods that use uncontrolled assumptions such as surrogate value functions and an infinite number of particles, SGDD does not rely on such approximations and eventually converges to the posterior distribution with robustness to score approximation and discretization error. Consequently, SGDD can be viewed as a natural and principled approach for posterior sampling using discrete diffusion models, leading to empirical benefits as we demonstrate in Section 4.
@@ 4.1 Experimental Setup
We evaluate our method across multiple domains, including synthetic data, DNA sequences, discrete image prior, and monophonic music, as listed in Table 1. We define the likelihood function for all inverse problems as $p({\mathbf{y}}|{\mathbf{x}})\ \propto\ \exp(-\|{\bm{G}}({\mathbf{x}})-{\mathbf{y}}\|/\sigma_{\mathbf{y}})$, while the likelihood functions for reward guiding tasks are $p({\mathbf{y}}|{\mathbf{x}})\ \propto\ \exp(\beta r({\mathbf{x}}))$.
Pretrained models. For synthetic data, we use a closed-form concrete score function. For real datasets, we train a SEDD [29] model with the uniform transition kernel (details in Appendix C).
Baselines. We compare our methods to existing approaches for discrete diffusion posterior sampling: DPS [7], SVDD-PM [27], and SMC [47]. The original version of DPS relies on gradient back-propagation through both the likelihood function and the pretrained diffusion model, making it unsuitable for discrete-domain posterior sampling. In our experiments, we consider its discrete analogy by replacing the gradient with a guidance rate matrix. SVDD-PM is originally designed to sample from a reward-guided distribution $p^{\beta}({\mathbf{x}})\ \propto\ p({\mathbf{x}})\exp(\beta r({\mathbf{x}}))$, but it is also adaptable to solving inverse problems. More implementation details are elaborated in Section C.2.
SGDD implementation details. We implement our method with a total of $K$ iterations. In each prior sampling step, we simulate the reverse continuous-time Markov chain with $H=20$ steps. Additional details on hyperparameter choices are provided in Section C.3.

### v3

@@ Abstract
We study the problem of posterior sampling in discrete-state spaces using discrete diffusion models. While posterior sampling methods for continuous diffusion models have achieved remarkable progress, analogous methods for discrete diffusion models remain challenging. In this work, we introduce a principled plug-and-play discrete diffusion posterior sampling algorithm based on split Gibbs sampling, which we call SGDD. Our algorithm enables reward-guided generation and solving inverse problems in discrete-state spaces. We demonstrate the convergence of SGDD to the target posterior distribution and verify this through controlled experiments on synthetic benchmarks. Our method enjoys state-of-the-art posterior sampling performance on a range of benchmarks for discrete data, including DNA sequence design, discrete image inverse problems, and music infilling, achieving more than $30\%$ improved performance compared to existing baselines. Our code is available at https://github.com/chuwd19/Split-Gibbs-Discrete-Diffusion-Posterior-Sampling.
@@ 1 Introduction
We validate SGDD on diverse inverse problems and reward-guided generation problems that involve discrete data. We demonstrate that SGDD converges to the true posterior distribution on synthetic data, that SGDD solves discrete inverse problems accurately and efficiently on discretized image data and monophonic music data, and that SGDD excels at reward-guided sampling on guided DNA generation. Our method demonstrates strong sampling quality in all experiments. For instance, we achieve a $42\%$ higher median activity in enhancer DNA design, an 8.36dB8.36\ $\mathrm${dB} improvement of PSNR values in solving the XOR inverse problem on MNIST, and over $2\times$ smaller Hellinger distance in music infilling, compared to existing methods.
@@ 2.2 Diffusion Posterior Sampling Methods
Existing diffusion posterior sampling methods leverage (continuous) diffusion models as plug-and-play priors and modify Eq.˜1 to generate samples from the posterior distribution $p({\mathbf{x}}|{\mathbf{y}})\ \propto\ p({\mathbf{x}})p({\mathbf{y}}|{\mathbf{x}})$. These methods can be broadly classified into four categories [9, 56].
Guidance-based methods [7, 38, 37, 24, 49, 55] perform posterior sampling by adding an additional likelihood score term $\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{y}}|{\mathbf{x}}_{t})$ to the score function in Eq.˜1, and modifies the reverse SDE as: $\mathrm{d}{\mathbf{x}}_{t}=-2\dot{\sigma}_{t}\sigma_{t}(\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{x}}_{t};\sigma_{t})+\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{y}}|{\mathbf{x}}_{t}))\mathrm{d}t+\sqrt{2\dot{\sigma}_{t}\sigma_{t}}\mathrm{d}{\mathbf{w}}_{t}.$ The intractable likelihood score term is approximated with various assumptions. However, these methods cannot be directly applied to discrete posterior sampling problems, where $\log p({\mathbf{y}}|{\mathbf{x}})$ is defined only on a finite support and thus does not have such gradient information.
Sequential Monte Carlo (SMC) methods sample batches of particles iteratively from a series of distributions, which converge to the posterior distribution in the limit. This class of methods [45, 12, 26, 50, 33, 34, 5] exploits the sequential diffusion sampling process to sample from the posterior distribution. Similar ideas have been applied in Li et al. [27] for reward-guided generation using discrete diffusion models, which rely on approximating a soft value function at each iteration.
@@
Variable splitting methods [51, 54, 57, 36, 52] decompose posterior sampling into two simpler alternating steps. A representative of this category is the split Gibbs sampler (see Section˜3), where the variable ${\mathbf{x}}_{t}$ is split into two random variables that both eventually converge to the posterior distribution. These methods are primarily designed for continuous diffusion models with Gaussian kernels, leveraging Langevin dynamics or optimization in Euclidean space as subroutines. Murata et al. [31] extends this category to VQ-diffusion models in the discrete latent space using the Gumbel-softmax reparameterization trick. However, this approach is limited to discrete latent spaces with an underlying continuous embedding due to its reliance on softmax dequantization. In this work, we generalize the split Gibbs sampler to posterior sampling in categorical data by further exploiting the unique properties of discrete diffusion models.
@@ 3.1 Split Gibbs for plug-and-play posterior sampling
where $D({\mathbf{x}},{\mathbf{z}};\eta)$ measures the distance between $\mathbf{x}$} and $\mathbf{z}$}, and $\eta>0$ is a parameter that controls the strength of regularization. While prior works [8, 51, 52] consider $D({\mathbf{x}},{\mathbf{z}};\eta)=\frac{\|{\mathbf{x}}-{\mathbf{z}}\|^{2}_{2}}{2\eta^{2}}$, the potential function can be generalized to any continuous function $D$, such that $D({\mathbf{x}},{\mathbf{z}};\eta)\to\infty$ as $\eta\to 0$ for any ${\mathbf{x}}\neq{\mathbf{z}}$. As shown in Section˜A.3, this ensures that both marginal distributions,
$\pi^{X}({\mathbf{x}};\eta):=\int\pi({\mathbf{x}},{\mathbf{z}};\eta)\mathrm{d}{\mathbf{z}},\ \ \text{and}\ \ \pi^{Z}({\mathbf{z}};\eta):=\int\pi({\mathbf{x}},{\mathbf{z}};\eta)\mathrm{d}{\mathbf{x}}$
converges to the posterior distribution $p({\mathbf{x}}|{\mathbf{y}})$. The decoupling of the prior and likelihood in Eq.˜7 enables split Gibbs sampling, which alternates between two steps:
@@ 2. Prior sampling step:
Prior work has studied split Gibbs posterior sampling for continuous diffusion models [8, 51, 52]. These methods specify the regularization potential as $D({\mathbf{x}},{\mathbf{z}};\eta)=\frac{\|{\mathbf{x}}-{\mathbf{z}}\|_{2}^{2}}{2\eta^{2}}$, which transforms Eq.˜9 into a Gaussian denoising problem solvable by a diffusion model. A key challenge, which we tackle, is developing a performant approach for the discrete diffusion setting that is easy to implement, and enjoys rigorous guarantees (i.e., sampling from the correct posterior distribution).
@@ 3.2 Prior Step with Discrete Diffusion Models
Suppose $p({\mathbf{x}})$ is a discrete-state distribution over $\mathcal{X}^{D}$ modeled by a diffusion process. We consider a discrete diffusion model with uniform transition kernel ${\bm{Q}}_{t}^{\mathsf{fw}}=\frac{1}{N}\mathbf{1}\mathbf{1}^{T}-{\bm{I}}$. To connect split Gibbs sampling to discrete diffusion models, we specify the potential function $D({\mathbf{x}},{\mathbf{z}};\eta)$ as:
@@
where $d({\mathbf{x}},{\mathbf{z}})$ denotes the Hamming distance between $\mathbf{x}$} and $\mathbf{z}$}. When $\eta\to 0^{+}$, the regularization potential $D({\mathbf{x}},{\mathbf{z}};\eta)$ goes to infinity unless $d({\mathbf{x}},{\mathbf{z}})=0$, ensuring the convergence of marginal distributions to $p({\mathbf{x}}|{\mathbf{y}})$. Given Eq.˜10, the prior step from Eq.˜9 can be written as
@@
where $\beta_{t}=\frac{N-1}{N}(1-e^{-\sigma_{t}})$. Thus, sampling from Eq.˜11 is equivalent to unconditional generation from $p({\mathbf{x}}_{0}|{\mathbf{x}}_{t})$ when $\tilde{\beta}=\beta_{t}$, i.e., $\eta=\sigma_{t}$. Therefore, we can solve the prior sampling problem by simulating a partial discrete diffusion sampler starting from $\sigma_{t}=\eta$ and ${\mathbf{x}}_{t}={\mathbf{z}}^{(k)}$.
@@ 3.3 Likelihood Sampling Step
With the potential function $D({\mathbf{x}},{\mathbf{z}};\eta)$ specified in Section˜3.2, at iteration $k$, the likelihood sampling step from Eq.˜8 can be written as:
@@
Since the unnormalized probability density function of $\pi({\mathbf{x}}={\mathbf{x}}^{(k)},{\mathbf{z}};\eta)$ is accessible, we can efficiently sample from Eq.˜13 using Markov Chain Monte Carlo methods.
@@
  Likelihood step following Eq.˜13: ${\mathbf{z}}^{(k)}\sim\pi({\mathbf{x}}={\mathbf{x}}^{(k)},{\mathbf{z}};\eta_{k}).$
  Prior step following Eq.˜11: ${\mathbf{x}}^{(k+1)}\sim\pi({\mathbf{x}},{\mathbf{z}}={\mathbf{z}}^{(k)};\eta_{k})$
@@ 3.4 Overall Algorithm
We now summarize the complete SGDD algorithm. Like in Wu et al. [51], our algorithm alternates between likelihood steps and prior steps while employing an annealing schedule $\{\eta_{k}\}$, which starts at a large $\eta_{0}$ and gradually decays to $\eta_{K}\to 0$. This annealing scheme accelerates the mixing time of the Markov chain, which we further provide some theoretical intuition in Section˜B.1; and ensures the convergence of $\pi^{X}({\mathbf{x}};\eta)$ and $\pi^{Z}({\mathbf{z}};\eta)$ to $p({\mathbf{x}}|{\mathbf{y}})$ as $\eta\to 0$. We present the complete pseudocode of our method in Algorithm˜1 and further explain the choice of the noise schedule in Section˜B.1.
@@ 3.5 Convergence of Split Gibbs Diffusion Posterior Sampling
[…]
where $\left\|\frac{s_{\theta}(\cdot;t)-s(\cdot;t)}{s(\cdot;t)}\right\|_{\infty}\leq\epsilon\leq 1$, and $L,M,c$ are positive constants defined in Appendix A.
We include the proof in Appendix˜A. Theorem 1 states that the average Fisher divergence of the non-stationary process to the stationary process in all prior steps converges at the rate of $O(1/K)$, up to a constant error term. This result extends our theoretical understanding of diffusion posterior sampling by generalizing the analysis [43, 51] on SDEs to general Markov processes using the free-energy-rate-functional-relative-Fisher-information (FIR) inequality [16]. Moreover, compared to existing analyses for continuous diffusion models, our analysis accounts not only for the imperfect score function but also for the discretization error in solving CTMCs as well.
Unlike some previous methods [27, 50] that rely on assumptions of surrogate value functions and an infinite number of particles, SGDD does not rely on such approximations and eventually converges to the posterior distribution with robustness to score approximation and discretization error. Consequently, SGDD can be viewed as a natural and principled approach for posterior sampling using discrete diffusion models, leading to empirical benefits as we demonstrate in Section˜4.
@@ 4.1 Experimental Setup
We evaluate our method across multiple domains, including synthetic data, DNA sequences, discrete image prior, and monophonic music, as listed in Table˜1. We define the likelihood function for all inverse problems as $p({\mathbf{y}}|{\mathbf{x}})\ \propto\ \exp(-\|{\bm{G}}({\mathbf{x}})-{\mathbf{y}}\|/\sigma_{\mathbf{y}})$, while the likelihood functions for reward guiding tasks are $p({\mathbf{y}}|{\mathbf{x}})\ \propto\ \exp(\beta r({\mathbf{x}}))$.
Pretrained models. For synthetic data, we use a closed-form concrete score function. For real datasets, we train a SEDD [29] model with the uniform transition kernel (details in Appendix˜C).
Baselines. We compare our methods to existing approaches for discrete diffusion posterior sampling: DPS [7], SVDD-PM [27], and SMC [50]. The original version of DPS relies on gradient back-propagation through both the likelihood function and the pretrained diffusion model, making it unsuitable for discrete-domain posterior sampling. In our experiments, we consider its discrete analogy by replacing the gradient with a guidance rate matrix. SVDD-PM is originally designed to sample from a reward-guided distribution $p^{\beta}({\mathbf{x}})\ \propto\ p({\mathbf{x}})\exp(\beta r({\mathbf{x}}))$, but it is also adaptable to solving inverse problems. More implementation details are elaborated in Section˜C.2.
SGDD implementation details. We implement our method with a total of $K$ iterations. In each prior sampling step, we simulate the reverse continuous-time Markov chain with $H=20$ steps. Additional details on hyperparameter choices are provided in Section˜C.3.